OpenStudy (anonymous):

plz check my answer What is the solution to the equation log 2 x – log 2 4 = 2 ? x = 1 x = 18 x = 8 --->My answer x = 16

OpenStudy (phi):

the problem is \[ \log_2(x) - \log_2(4) = 2\] use the property \[ \log(a)- \log(b) = \log(a/b) \] to combine the logs. can you do that ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

you get \[ \log_2\left(\frac{x}{4}\right) = 2\]

OpenStudy (anonymous):

so it wou ld be 4x=2 then i get x=2 but that isnt one of the choices

OpenStudy (phi):

you get \[ \log_2\left(\frac{x}{4}\right) = 2 \] to get rid of a log, make it the exponent of its *base* make both sides have a base of 2: \[ 2^{\log_2\left(\frac{x}{4}\right)}= 2^2\] the mess on the left simplifies to x/4 (which is why we do this) \[ \frac{x}{4}= 2^2\] can you finish?

OpenStudy (anonymous):

yes!!! x/4=4 then i divide 4 by 4 and i get 1?

OpenStudy (phi):

first step is ok \[ \frac{x}{4} = 4\] what is x ? if x is 1, you would get \[ \frac{1}{4} = 4\] which is not correct. 1/4 does not equal 4 so x is not 1. we could try x=2, and x=3 and so on, but there are an awful lot of numbers to check. one way to find x is multiply both sides by 4 \[ 4\cdot \frac{x}{4} = 4 \cdot 4\] on the left side, 4 "up top" and 4 "below" means 4/4 or 4 divided by 4, or just 1 you get \[ x= 4 \cdot 4\]

OpenStudy (anonymous):

oh ok i see thanks so much for helping me

OpenStudy (anonymous):

thanks i will watch it